# Sven Lilge and Timothy D. Barfoot Incorporating Control Inputs in Continuous-Time Gaussian Process State Estimation for Robotics

### v2

@@ Abstract
Continuous-time batch state estimation using Gaussian processes is an efficient approach to estimate the trajectories of robots over time. In the past, relatively simple physics-motivated priors have been considered for such approaches, using assumptions such as constant velocity or acceleration. This paper presents an approach to incorporating exogenous control inputs, such as velocity or acceleration commands, into the continuous Gaussian process state-estimation framework. It is shown that this approach generalizes across different domains in robotics, making it applicable to both the estimation of continuous-time trajectories for mobile robots and the estimation of quasi-static continuum robot shapes. Results show that incorporating control inputs leads to more informed priors, potentially requiring less measurements and estimation nodes to obtain accurate estimates. This makes the approach particularly useful in situations in which limited sensing is available.
1 Introduction
@@
Throughout this manuscript, we are specifically focusing on continuous-time batch state estimation using Gaussian processes (GPs) to represent the robot’s state and uncertainty continuously over time [2]. This approach involves formulating a suitable prior distribution of continuous robot trajectories. Current formulations are usually based on simple physics models, making specific assumptions about the motion of the robot. In [3] a white-noise-on-acceleration (WNOA) prior is employed, assuming robot motion with constant velocity. This approach has been extended to a white-noise-on-jerk (WNOJ) prior in [4], which assumes constant acceleration. These formulations have been shown to also be applicable to the continuous estimation of the shape of continuum robots [5, 6, 7], substituting time with arclength and velocity with strain. Fig. 1 highlights this analogy. For mobile robots, we seek to estimate their pose and velocity along a continuous trajectory, where the continuous variable is time. For continuum robots, an analogous estimation problem can be formulated, in which we seek to estimate the robot’s pose and strain along their continuous shape. The continuous variable is now arclength and instead of estimating a trajectory over time, we are estimating a quasi-static shape in space. In both cases, the robot’s pose and their derivative are estimated along a continuous smooth curve (trajectory or shape). In many cases, the GP prior formulations represent the mobile robot motions or continuum robot shapes well. However, besides the assumptions of constant velocity (or strain) or acceleration (equivalent to the strain derivative for continuum robots), no additional knowledge about the robot state is incorporated into the prior, such as exogenous control inputs to the system. For mobile robots, these inputs could be known velocity or acceleration commands, while for continuum robots they could consist of known actuation forces and moments acting on their structure. Incorporating such information directly into the GP formulation can lead to more informed prior distributions, representing the mobile robot trajectory or continuum robot shape more accurately. Ultimately, this can mean that less measurements are required to accurately estimate the continuous robot state, when compared to the traditional prior formulations.
@@ 1.1 Related Work
Initial approaches to infer the continuous state between two discrete states of a batch estimation approach relied on simple linear interpolation [8, 9, 10], which is a rather restrictive assumption about the robot’s motion. Following this, first parametric approaches were proposed, using temporal basis functions to represent continuous and smooth robot trajectories [1, 11]. A nonparametric approach, relying only on time as its sole parameter, was proposed in [2, 12], using GPs to represent the continuous state in 2D and 3D. Both of these works allow for a variety of different GP priors featuring relatively large and dense kernel matrices, ultimately limiting computational efficiency. A more efficient GP prior was proposed in [13], which is derived from linear, time-varying stochastic differential equations driven by white noise, leading to an exactly sparse kernel matrix by exploiting the Markov property. The proposed prior employs white noise on acceleration for 2D robot trajectories, leading to a constant velocity assumption. The WNOA prior was extended to in [3], while [4] proposed an extension to consider white-noise-on-jerk prior formulations, assuming a constant acceleration instead of constant velocities. Both the WNOA and WNOJ priors are relatively simple physically motivated formulations that do not consider exogenous control inputs. Instead, velocity or acceleration inputs are often incorporated as measurements of the state [14, 15]. The continuous-time motion prior for the estimation of mobile robot trajectories was further repurposed to the continuous estimation of continuum robot shapes. In [5] the WNOA prior was adopted to estimate the quasi-static continuous shape and strain of continuum robots, considering arclength as the continuous parameter instead of time. This approach was extended to coupled robot topologies in [7], while [6] adopts a WNOJ prior to additionally estimate continuously applied forces to the continuum body. A recent study compares the GP-based continuous-time estimation approach with parametric spline-based methods, noting that results are similar regarding the achieved accuracies and computation times [16].
@@
Another approach to continuous state estimation using GPs is presented in [17] and [18]. Here, GPs are used to pre-integrate high-rate IMU measurements, creating pseudo-measurements between two discretely estimated states. The idea of pre-integrating high-rate measurements was first discussed in [19], which uses numerical integration schemes, and is the backbone of many discrete-time batch estimation schemes [20, 21]. Contrary to these traditional methods, the approach in [17] and [18] uses a continuous GP latent space representation fitted to measurements, which is then integrated in closed form (i.e., Bayesian quadrature). This results in a smooth pseudo-measurement, which is then used as a factor between two discrete states in a batch estimation approach. Unlike the continuous-time estimation methods discussed before, this work does not feature an underlying continuous motion prior and carries out estimation purely in discrete time after pre-integration. It is thus unable to estimate the continuous mean and covariance of a robot’s trajectory and could fail if the high-rate measurements between two discrete states drop out.
@@ 1.2 Contributions
As previously stated, this work presents a direct extension of the established WNOA Gaussian process prior [3, 5] to include exogenous control inputs. The derivations presented throughout this manuscript allow for the inclusion of inputs on both the velocity and acceleration. To the best of the authors’ knowledge, this work presents the first GP-based continuous-time/shape estimation approach that combines a physically motivated motion prior with such control inputs. By combining ideas from the pre-integration literature [17, 18] with previously proposed continuous-time motion priors [3, 5], a single factor between two discrete robot states is derived, which can then be used in a batch estimation approach. While this is similar to what is currently done in pre-integration approaches, we are able to interpolate the continuous robot state over time, considering both the motion model and inputs. This allows us to recover the local shape and uncertainty of the trajectory between two discrete states at any query time. Additionally, in the case in which no known inputs are available, the proposed method gracefully falls back to the underlying WNOA motion prior for estimation and interpolation.
Throughout this manuscript, we will derive formulations that consider continuous control inputs in , allowing us to approximate the resulting GP prior in closed form. The proposed approach is applied to both the estimation of continuous mobile robot trajectories and continuum robot shapes, highlighting advantages over the existing GP motion priors. We show that considering control inputs is particularly useful when the available measurement data is sparse.
@@ 2.4 Gaussian Process Interpolation
Lastly, we are interested in interpolating the GP posterior at any arbitrary time along the continuous trajectory. For this, we start with the general interpolation equations for the GP mean stated in [22, Sec. 11.3]:
@@
Now, we can solve for and considering the mapping between global and local pose variables in (2) and (3). The covariance of the posterior GP can be interpolated similarly. Derivations are omitted for the sake of space, but details can be found in [22, Sec. 11.3] and [23]. This interpolation scheme can further be used during the MAP optimization to incorporate measurement terms at query times different from our discrete times [15].
@@ 3.1 Dataset
We use a mobile robot dataset similar to the one reported in [2]. In our dataset, a mobile robot is driving within an indoor, planar environment featuring 17 tubular landmarks. The overall robot trajectory includes 21 minutes of driving. The robot is equipped with a laser rangefinder, measuring the distance to each landmark at 10 Hz. Using wheel odometry readings, measurements of the robot’s body velocity are available at 10 Hz as well. Ground-truth for the robot trajectory is provided by a Vicon motion capture system.
@@ 3.2 Localization Evaluation
For evaluation, we consider the localization problem, where landmark locations are known. We are using the wheel odometry readings as piecewise-linear velocity inputs at a rate of 10 Hz to create a GP motion prior. Landmark distance readings are included as discrete measurements, which are fused with the prior to create a posterior estimate. We employ a MAP estimation scheme that computes the posterior distribution of the entire trajectory in a single batch. We are particularly interested in how the proposed approach performs compared to the conventional WNOA GP prior baseline, in which odometry readings are considered as velocity measurements of the state, instead of as inputs. For both methods, we optimize the hyperparameters of the prior, namely , on a subset of the available mobile robot driving data. For the WNOA method, we additionally optimize the covariance of the odometry velocity measurements. The covariance for the landmark distance measurements is given.
@@
| Prior Without Inputs | Prior Without Inputs (Exclude states w/o dist. meas.) | Prior With Inputs (Ours) | Prior With Inputs (Ours) (Exclude states w/o dist. meas.) | |||||||||
| in s | Pos. RMSE in m | Rot. RMSE in rad | Time in s | Pos. RMSE in m | Rot. RMSE in rad | Time in s | Pos. RMSE in m | Rot. RMSE in rad | Time in s | Pos. RMSE in m | Rot. RMSE in rad | Time in s |
| 0.5 | 0.0244 | 0.085 | 12.9 | 0.0248 | 0.0846 | 3.2 | 0.0259 | 0.0827 | 21.9 | 0.0259 | 0.0827 | 3.1 |
| 1 | 0.0249 | 0.0855 | 10.6 | 0.0284 | 0.0872 | 2.3 | 0.0261 | 0.0825 | 21.1 | 0.0263 | 0.0825 | 2.5 |
| 2 | 0.027 | 0.0871 | 10.8 | 0.0424 | 0.0996 | 2.3 | 0.0276 | 0.0833 | 18.3 | 0.0285 | 0.0833 | 2.8 |
| 3 | 0.0293 | 0.0872 | 10.6 | 0.0693 | 0.1134 | 2 | 0.0299 | 0.0827 | 20.8 | 0.0317 | 0.083 | 3.7 |
| 4 | 0.032 | 0.0878 | 9.1 | 0.0953 | 0.1403 | 1.8 | 0.031 | 0.083 | 19.6 | 0.0353 | 0.0836 | 4.5 |
| 5 | 0.0348 | 0.088 | 9.1 | 0.1309 | 0.1633 | 1.8 | 0.0309 | 0.0829 | 21.1 | 0.0368 | 0.0837 | 6 |
| 6 | 0.0358 | 0.0887 | 10.8 | 0.1706 | 0.1904 | 1.7 | 0.0347 | 0.0837 | 23.9 | 0.0464 | 0.0853 | 7.8 |
| 7 | 0.0413 | 0.0907 | 10.4 | 0.2296 | 0.2281 | 2.2 | 0.0371 | 0.0849 | 20.8 | 0.0585 | 0.0856 | 10.1 |

### v3

@@ Abstract
Continuous-time batch state estimation using Gaussian processes is an efficient approach to estimate the trajectories of robots over time. In the past, relatively simple physics-motivated priors have been considered for such approaches, using assumptions such as constant velocity or acceleration. This paper presents an approach to incorporating exogenous control inputs, such as velocity or acceleration commands, into the continuous Gaussian process state-estimation framework. It is shown that this approach generalizes across different domains in robotics, making it applicable to both the estimation of continuous-time trajectories for mobile robots and the estimation of quasi-static continuum robot shapes. Results show that incorporating control inputs leads to more informed priors, potentially requiring less measurements and estimation nodes to obtain accurate estimates. This makes the approach particularly useful in situations in which limited sensing is available. For example, in a mobile robot localization experiment with sparse landmark distance measurements and frequent odometry control inputs, our approach provides accurate trajectory estimates with root-mean-square errors around 3-4 cm and 4-5 degrees, even with time intervals up to five seconds between discrete estimation nodes, which significantly reduces computation time.
keywords:
Continuous-Time State Estimation, Gaussian Process Regression, Robot Localization, Pose Estimation and Registration, Group-Theoretic Methods1 Introduction
@@
Throughout this manuscript, we are specifically focusing on continuous-time batch state estimation using Gaussian processes (GPs) to represent the robot’s state and uncertainty continuously over time [33]. This approach involves formulating a suitable prior distribution of continuous robot trajectories. Current formulations are usually based on simple physics models, making specific assumptions about the motion of the robot. In [1] a white-noise-on-acceleration (WNOA) prior is employed, assuming robot motion with constant velocity. This approach has been extended to a white-noise-on-jerk (WNOJ) prior in [30], which assumes constant acceleration. These formulations have been shown to also be applicable to the continuous estimation of the shape of continuum robots [22, 13, 23], substituting time with arclength and velocity with strain. Fig. 1 highlights this analogy.
For mobile robots, we seek to estimate their pose and velocity along a continuous trajectory, where the continuous variable is time. For continuum robots, an analogous estimation problem can be formulated, in which we seek to estimate the robot’s pose and strain along their continuous shape. The continuous variable is now arclength and instead of estimating a trajectory over time, we are estimating a quasi-static shape in space. In both cases, the robot’s pose and its derivative are estimated along a continuous, smooth curve – representing either a trajectory or shape – by fusing the Gaussian process prior with noisy state observations of various types. For instance, in mobile robotics, noisy landmark observations are commonly used to relate the robot’s pose and orientation to landmarks in the environment. In continuum robotics, electromagnetic tracking sensors might provide noisy direct measurements of the pose.
In many cases, the GP prior formulations represent the mobile robot motions or continuum robot shapes well. However, besides the assumptions of constant velocity (or strain) or acceleration (equivalent to the strain derivative for continuum robots), no additional knowledge about the robot state is incorporated into the prior, such as exogenous control inputs to the system. For mobile robots, these inputs could be known velocity or acceleration commands, while for continuum robots they could consist of known actuation forces and moments acting on their structure. Incorporating such information directly into the GP formulation can lead to more informed prior distributions, representing the mobile robot trajectory or continuum robot shape more accurately. Ultimately, this can mean that less measurements are required to accurately estimate the continuous robot state, when compared to the traditional prior formulations.
@@ 1.1 Related Work
In the following, we review commonly used continuous-time state estimation methods. For a more comprehensive introduction, we refer interested readers to a recent survey on these approaches [29].
Initial approaches to infer the continuous state between two discrete states of a batch estimation approach relied on simple linear interpolation. For instance, the works in [6, 7] assume constant velocities over short time periods to linearly interpolate the robot’s pose for LiDAR Simultaneous Localization and Mapping (SLAM). Similarly, [35, 17] use linear pose interpolations to correct for motion distortion during LiDAR odometry. While linear interpolation schemes are easy to implement and fast to compute, their accuracy is generally limited, as they do not necessarily correspond to the actual robot motion.
As an alternative, first parametric continuous-time state estimation approaches were proposed, using temporal basis functions to represent continuous and smooth robot trajectories [16]. While initial investigations were limited to deal with states expressed in vector spaces, this approach has since been extended to Lie group formulations [24, 26]. Recent advancements mostly focus on increasing computational efficiency of these approaches, e.g., by deriving analytical Jacobians for Lie group formulations [28, 31, 21].
@@
A nonparametric approach, relying only on time as its sole parameter, was proposed in [33, 32], using GPs to represent the continuous state in 2D and 3D. Both of these works allow for a variety of different GP priors featuring relatively large and dense kernel matrices, ultimately limiting computational efficiency. A more efficient GP prior was proposed in [4], which is derived from linear, time-varying stochastic differential equations driven by white noise, leading to an exactly sparse inverse kernel matrix by exploiting the Markov property. The proposed prior employs white noise on acceleration for 2D robot trajectories, leading to a constant velocity assumption. The WNOA prior was extended to in [1], while [30] proposed an extension to consider white-noise-on-jerk prior formulations, assuming a constant acceleration instead of constant velocities. In [34] the Singer prior is proposed, which introduces an additional parameter to the GP prior, representing latent accelerations. Learning this parameter from data can potentially lead to priors that represent the robot’s motion more accurately than the WNOA and WNOJ priors. Lastly, derivations for formulations of the WNOA GP prior for general matrix Lie groups, beyond just , are proposed in [11, 12]. All of the aforementioned GP priors are relatively simple physically motivated formulations that do not consider exogenous control inputs. Instead, velocity or acceleration inputs are often incorporated as measurements of the state [9, 8].
The continuous-time motion prior for the estimation of mobile robot trajectories was further repurposed to the continuous estimation of continuum robot shapes. In [22] the WNOA prior was adopted to estimate the quasi-static continuous shape and strain of continuum robots, considering arclength as the continuous parameter instead of time. This approach was extended to coupled robot topologies in [23], while [13] adopts a WNOJ prior to additionally estimate continuously applied forces to the continuum body. A recent study compares the GP-based continuous-time estimation approach with parametric spline-based methods, noting that results are similar regarding the achieved accuracies and computation times [18].
@@
Another approach to continuous state estimation using GPs is presented in [20] and [19]. Here, GPs are used to pre-integrate high-rate IMU measurements, creating pseudo-measurements between two discretely estimated states. The idea of pre-integrating high-rate measurements was first discussed in [25], which uses numerical integration schemes, and is the backbone of many discrete-time batch estimation schemes [14, 15]. Contrary to these traditional methods, the approach in [20] and [19] uses a continuous GP latent space representation fitted to measurements, which is then integrated in closed form (i.e., Bayesian quadrature). This results in a smooth pseudo-measurement, which is then used as a factor between two discrete states in a batch estimation approach. Unlike the continuous-time estimation methods discussed before, this work does not feature an underlying continuous motion prior and carries out estimation purely in discrete time after pre-integration. It is thus unable to estimate the continuous mean and covariance of a robot’s trajectory and could fail if the high-rate measurements between two discrete states drop out. A similar pre-integration approach is presented in [10]. Here, conventional on-manifold pre-integration [15] is combined with a data-driven dynamics model, to incorporate dynamic motion model factors into the discrete-time batch estimation of drone trajectories.
To conclude, the main limitations of current approaches lie in the simplicity of the physical assumptions underlying most continuous-time state estimation methods, such as constant velocity or constant acceleration. While these assumptions work well if the robot motion is close to these conditions, they fail to adequately represent more complex or varying motion. Pre-integration methods partially offer a solution by capturing robot motion through pre-integrated factors considering high-rate measurements. However, they revert back to a discrete-time estimation problem at runtime, sacrificing the ability to continuously query the state at any desired time.
Consequently, a key question remains open: how can more intricate physics models, especially ones incorporating exogenous control inputs, be integrated into continuous-time state estimation frameworks? Addressing this challenge is crucial for formulating priors that more accurately represent the actual motion of robots. This issue is also identified as an open research question in a recent survey on continuous-time state estimation methods [29]. Tackling this problem has the potential to enhance both the accuracy and applicability of continuous-time state estimation methods.
@@ 1.2 Contributions
This work presents a direct extension of the established WNOA Gaussian process prior [1, 22] to include exogenous control inputs. The derivations presented throughout this manuscript allow for the inclusion of inputs on both the velocity and acceleration. To the best of the authors’ knowledge, this work presents the first GP-based continuous-time estimation approach that combines a physically motivated motion prior with such control inputs. By combining ideas from the pre-integration literature [20, 19] with previously proposed continuous-time motion priors [1, 22], a single factor between two discrete robot states is derived, which can then be used in a batch estimation approach.
While our method shares similarities with existing approaches in the current literature, there are some striking differences that set it apart. As discussed in the previous section, continuous-time state estimation methods for robotics are traditionally based on temporal basis functions, such as splines, or GPs [29, 18]. While spline formulations are not directly tied to motion process models, GPs are motivated by the underlying physics of robot motion, theoretically allowing the inclusion of complex process models and control inputs. However, previous GP-based methods have been limited to simple motion models such as constant velocity or constant acceleration [1, 30].
Our approach makes contributions to the field by introducing a GP prior capable of representing more complex robot motions defined by exogenous control inputs. This is the first time such an extension has been proposed. Consequently, our priors are more informed, more general, and potentially more accurate. Similar to pre-integration methods [20, 19], which pre-integrate IMU measurements to compute relative pseudo-measurement factors, our motion prior factors describe the relative motion between discrete states in time. However, unlike pre-integration methods, which estimate the states at discrete times and are unable to query it at arbitrary times after computation of the factors, our GP-based approach allows straightforward interpolation of the mean and covariance of the robot’s state at any query time. This enables the recovery of the local shape and uncertainty of the trajectory between discrete states. Additionally, in the case in which no known inputs are available, the proposed method gracefully falls back to the underlying WNOA motion prior for estimation and interpolation.
In conclusion, this work advances continuous-time state estimation, introducing a GP-based framework that incorporates process models considering exogenous control inputs, potentially paving the way for more versatile continuous-time estimation approaches in robotics. Throughout this manuscript, we derive formulations that consider continuous control inputs in , allowing us to approximate the resulting GP prior in closed form. The proposed approach is applied to both the estimation of continuous mobile robot trajectories and quasistatic continuum robot shapes, highlighting advantages over the existing GP motion priors. We show that considering control inputs is particularly useful when the available measurement data is sparse.
@@ 2.4 Gaussian Process Interpolation
Lastly, we are interested in interpolating the GP posterior at any arbitrary time along the continuous trajectory. For this, we start with the general interpolation equations for the GP mean stated in [3, Sec. 11.3]:
@@
Now, we can solve for and considering the mapping between global and local pose variables in (2) and (3). The covariance of the posterior GP can be interpolated similarly. Derivations are omitted for the sake of space, but details can be found in [3, Sec. 11.3] and [2]. This interpolation scheme can further be used during the MAP optimization to incorporate measurement terms at query times different from our discrete times [8].
@@ 3.1 Dataset
We use a mobile robot dataset similar to the one reported in [33], collected in an indoor, planar environment featuring 17 tubular landmarks. Both the robot and the tubes were equipped with reflective markers tracked by a ten-camera Vicon motion capture system. This system provides sub-millimeter accuracy for the position of each reflective marker, which we use as ground truth for the robot’s trajectory. It further allows us to determine the exact location of the tubular landmarks, which we assume to be known during our localization experiment.
The mobile robot was driven around for 21 minutes. The dataset includes laser rangefinder scans from a Hokuyo URG-04LX sensor, consisting of 681 range measurements spread over a 240∘ horizontal field of view centered straight ahead, logged at 10 Hz. Additionally, wheel odometry readings were recorded at 10 Hz to provide estimates of the robot’s body-centric velocity. The experimental setup and ground-truth data of the recorded trajectory are shown in Fig. 4. It can be seen that the robot’s path is quite twisted and looping.
@@ 3.2 Localization Evaluation
For evaluation, we consider the localization problem, where landmark locations are known. We are using the wheel odometry readings as piecewise-linear velocity inputs at a rate of 10 Hz to create a GP motion prior. Landmark distance readings are included as discrete measurements, which are fused with the prior to create a posterior estimate. We employ a MAP estimation scheme that computes the posterior distribution of the entire trajectory in a single batch. We are particularly interested in how the proposed approach performs compared to the conventional WNOA GP prior baseline, in which odometry readings are considered as velocity measurements of the state, instead of as inputs. We believe that this comparison is meaningful, as the WNOA prior is one of the most widely used and well-established methods for continuous-time state estimation [29], which our proposed method directly extends. While other methods, such as spline-based techniques, exist, they are fundamentally different from the GP-based methods considered in this work, with detailed comparisons reported in [18]. For both our proposed method and the chosen baseline, we optimize the hyperparameters of the prior, namely , on a subset of the available mobile robot driving data. For the WNOA method, we additionally optimize the covariance of the odometry velocity measurements. The covariance for the landmark distance measurements is given. The resulting hyperparameters used throughout the experiment are summarized in Table 1. Here, the prior covariances include one translational and one rotational degree of freedom, considering a non-holonomic mobile robot in a two-dimensional environment. Since our derivations are formulated in , the remaining degrees of freedom are locked during optimization. Similarly, the velocity measurement covariance includes two values, as we measure the forward and yaw velocities of the mobile robot using wheel odometry readings.
| Prior Covariance (Prior with Inputs) | Landmark Distance Measurement Covariance |
| Prior Covariance (Prior without Inputs) | Velocity Measurement Covariance |
@@
| Prior Without Inputs | Prior With Inputs (Ours) | |||||||||
| in s | Pos. Error (cm) | Rot. Error (°) | Time in s | Pos. Error (cm) | Rot. Error (°) | Time in s | ||||
| RMSE | Max | RMSE | Max | RMSE | Max | RMSE | Max | |||
| 0.5 | 2.44 | 16.37 | 4.87 | 15.15 | 12.9 | 2.59 | 15.71 | 4.74 | 15.14 | 21.9 |
| 1 | 2.49 | 17.29 | 4.90 | 15.12 | 10.6 | 2.61 | 16.10 | 4.73 | 15.15 | 21.1 |
| 2 | 2.70 | 17.98 | 4.99 | 15.34 | 10.8 | 2.76 | 16.72 | 4.77 | 15.17 | 18.3 |
| 3 | 2.93 | 17.44 | 5.00 | 15.58 | 10.6 | 2.99 | 15.99 | 4.74 | 15.70 | 20.8 |
| 4 | 3.20 | 18.36 | 5.03 | 15.96 | 9.1 | 3.10 | 17.10 | 4.76 | 15.68 | 19.6 |
| 5 | 3.48 | 19.54 | 5.04 | 16.88 | 9.1 | 3.09 | 17.82 | 4.75 | 17.10 | 21.1 |
| 6 | 3.58 | 18.45 | 5.08 | 16.06 | 10.8 | 3.47 | 17.19 | 4.80 | 16.24 | 23.9 |
| 7 | 4.13 | 24.87 | 5.20 | 16.59 | 10.4 | 3.71 | 25.00 | 4.86 | 17.02 | 20.8 |
| Prior Without Inputs (Exclude states w/o dist. meas.) | Prior With Inputs (Ours) (Exclude states w/o dist. meas.) | |||||||||
| in s | Pos. Error (cm) | Rot. Error (°) | Time in s | Pos. Error (cm) | Rot. Error (°) | Time in s | ||||
| RMSE | Max | RMSE | Max | RMSE | Max | RMSE | Max | |||
| 0.5 | 2.48 | 15.96 | 4.85 | 14.63 | 3.2 | 2.59 | 15.77 | 4.74 | 15.14 | 3.1 |
| 1 | 2.84 | 17.09 | 5.00 | 15.57 | 2.3 | 2.63 | 16.35 | 4.73 | 15.18 | 2.5 |
| 2 | 4.24 | 13.80 | 5.71 | 19.16 | 2.3 | 2.85 | 17.33 | 4.77 | 15.32 | 2.8 |
| 3 | 6.95 | 27.17 | 6.50 | 25.88 | 2 | 3.17 | 16.45 | 4.76 | 15.88 | 3.7 |
| 4 | 9.56 | 32.04 | 8.07 | 33.97 | 1.8 | 3.53 | 17.86 | 4.79 | 16.12 | 4.5 |
| 5 | 13.08 | 47.20 | 9.35 | 41.73 | 1.8 | 3.68 | 17.83 | 4.80 | 17.48 | 6 |
| 6 | 17.09 | 60.60 | 10.93 | 45.66 | 1.7 | 4.64 | 16.93 | 4.89 | 16.49 | 7.8 |
| 7 | 22.83 | 71.80 | 13.10 | 52.46 | 2.2 | 5.85 | 23.55 | 4.90 | 16.58 | 10.1 |